### v2

# New Construction of Locally Perfect Nonlinear Functions with Application to Sequences Sets with Low Ambiguity Zone

Abstract
Low Ambiguity Zone (LAZ) sequences play a pivotal role in modern integrated sensing and communication (ISAC) systems. Recently, Wang et al.[1] proposed a definition of locally perfect nonlinear functions (LPNFs) and constructed three classes of both periodic and aperiodic LAZ sequence sets with flexible parameters by applying such functions and interleaving method. Some of these LAZ sequence sets are asymptotically optimal with respect to the Ye-Zhou-Liu-Fan-Lei-Tang bounds under certain conditions. In this paper, we proceed with the constructions of the three new classes of LPNFs with new parameters. By using these LPNFs, we also present a series of LAZ sequence sets with more flexible parameters, addressing the limitations of existing parameter choices. Furthermore, our results show that one of these classes is asymptotically optimal in both the periodic and aperiodic cases, respectively.
Index Terms
Low Ambiguity Zone (LAZ), Locally Perfect Nonlinear Function (LPNF), LAZ Sequence set.
I Introduction
To design sequence sets with various capabilities is one of the significant topics in the research field of communication and radar systems. A critical challenge now lies in the design of sequences that exhibit desirable correlation properties not only in the time domain but also across phase shifts, as effectively captured by the ambiguity function (AF). The AF is a valuable criterion for evaluating the communication and sensing capabilities of sequences[2], sequences exhibiting zero or low ambiguity function sidelobes are essential for mitigating the Doppler effect in advanced Integrated Sensing and Communication (ISAC) systems, enabling accurate signal detection and processing in highly dynamic environments[3]. In 2013, Ding and Feng et al.[4] investigated the theoretical lower bound of the maximum sidelobe of the ambiguity function (AF) based on the Welch bound. Additionally, they presented sequences possessing favorable AF characteristics. Notably, a single sequence among these achieves the desired theoretical lower bound across the entire delay-Doppler (DD) region. Moreover, Wang and Gong et al.[5, 6, 7] proposed several classes of sequences and sequence sets with low AF magnitudes in the entire DD region, leveraging the excellent algebraic properties of additive and multiplicative characters over finite fields. In addition to constructing sequences with low AF sidelobes based on finite field theory, researchers have also attempted to design sequences through optimization algorithms[8, 9, 10, 11]. However, some of these sequences exhibit low AF sidelobes only in specific DD regions.
In fact, for many practically interesting applications in ISAC system, the Doppler frequency range is typically much narrower than the bandwidth of the transmitted signal. Therefore, it’s likely that there’s no need to take into account the entire duration of the signal. In other words, in the study of different applications, it is usually sufficient to focus on the AF sidelobes in the region relevant to specific application requirements. Based on this idea, in the past five years, people have made breakthroughs in the research of AF. In reference [12], Ye et al. first introduced a new definition regarding the low AF sidelobes in a specific region, termed the low/zero ambiguity zone (LAZ/ZAZ). Therefore, they referred to the sequence sets that meet the requirements of the LAZ/ZAZ as LAZ/ZAZ sequence sets, and derived the theoretical lower bounds for the periodic and aperiodic AF sidelobes of unimodular sequence sets in the LAZ, respectively. These bounds are referred to as “Ye-Zhou-Liu-Fan-Lei-Tang bounds” in reference [1], for brevity, we use the name “YZLFLT bounds” in this paper. By applying cubic sequences, certain quadratic phase sequences and cyclic difference set, they also presented four types of the most suitable constructions of sequence family, and these sequences achieve the lower bounds. Cao et al.[13] subsequently identified that a binary sequence set proposed in [14] possesses low ambiguity characteristics over a large area. Note that these sequences are interleaved sequences. Thereafter, Tian et al.[15] defined a set of ZAZ/LAZ sequences based on the perfect nonlinear functions (PNFs) and a new class of mapping functions. Very recently, Wang et al.[1] generalized the concept of PNFs, introducing locally perfect nonlinear functions (LPNFs) and presented a construction of LPNFs. And then, by directly applying these LPNFs and interleaving method, they gave a series of LAZ sequence sets with flexible parameters. Such sequence sets with selecting appropriate parameters are asymptotically optimal with respect to YZLFLT bounds both periodic and aperiodic cases, respectively. Consequently, this drives us to develop sequences with zero or low AF sidelobes characteristics in desired regions. The specific parameters of the above mentioned sequences are listed in Table I. In column 6 of Table I, P.O, P.AO, and AP.AO mean optimal with periodic, asymptotically optimal with periodic, and asymptotically optimal with aperiodic, respectively.
TABLE I: Asymptotically Optimal or Optimal LAZ/ZAZ sequence sets
  Ref. Length Set size $Z_{x}$ $Z_{y}$ Optimality   Constraints
  [4] $q-1$ $1$ $q-1$ $q-1$ P.O. $q=p^{l}$, $p$ is a prime.
[12] $p$ $1$ $p$ $p$ P.O. $p$ is an odd prime.
[12] $L$ $1$ $\frac{L}{r}$ $r$ P.O. $\gcd(a,L)=1\,\textrm{if $L$ is odd },$ $r=\gcd(2a,L)>1$.
[12] $L$ $N$ $\frac{\lfloor L/N\rfloor}{r}$ $r$ P.O, if $N|L$. $\gcd(a,L)=1\,\textrm{if L is odd },$ $r=\gcd(2a,L)>1$.
[15] $MN^{2}$ $MN$ $\lfloor\frac{N}{K}\rfloor$ $K$ P.AO. $K<N$, $\gcd(K,N)=1$.
[15] $N(KN+P)$ $N$ $N$ $K$ P.AO. $\gcd(P,NK)=1$.
[15] $p(p-1)$ $p$ $p-1$ $p$ P.AO. $p$ is an odd prime.
[1] $N^{2}$ $N$ $p$ $N$ P.AO, AP.AO. $N$  is an odd number, $p|N(\textrm{the smallest prime})$.
[1] $NK$ $N$ $p$ $K-N+1$ P.AO, AP.AO. $N$  is an odd number, $N<K<2N-1$, $p|N(\textrm{the smallest prime})$.
Theorem5 (i) $NK$ $N$ $p-1$ $K$ P.AO, AP.AO. $K=p^{e},\textrm{$p$\,is an odd prime},$ $N=\varphi(p^{e})$.
Theorem5 (ii) $NK$ $N$ $N$ $p$ P.AO $K=p^{e},\textrm{$p$\,is an odd prime},$ $N=\varphi(p^{e})$.
Theorem6 $NK_{1}$ $N$ $p-1$ $K_{1}-K+2$ P.AO, AP.AO. $K=p^{e},\textrm{$p$\,is an odd prime},$ $K<K_{1}<2K$, $N=\varphi(p^{e})$.
In this paper, we show some generic new constructions on LPNFs in Section 3, and by using these new LPNFs, we present a series of LAZ sequence sets with new parameters. Some sequence sets of the proposed are asymptotically optimal with respect to YZLFLT bounds both periodic and aperiodic cases, respectively. Before doing our main results, in Section 2, we introduce some preliminaries on LAZ sequence sets and LPNFs, including some related results we need in this paper.
II LAZ Sequence Sets and Locally Perfect Nonlinear Functions
In this section we recall some essential concepts concerning LAZ sequence sets, locally perfect nonlinear functions (LPNFs), as well as results that are needed.
II-A LAZ Sequence Sets
In what follows, let $\mathbb{C}$ denote the set of all complex numbers and $\mathcal{S}$ denote the set of sequences of length $D$ which has a size of $M$, all sequences $\mathbf{a}=(a(0),a(1),a(2),\cdots,a(D-1))$ are regarded as unimodular sequences with length $D$. Namely, ${a}(k)\in\mathbb{C}$ and $|{a}(k)|=1$ for all $0\leq k\leq{D-1}$. In particular, $D$ is the period of $\mathbf{a}$ if ${a}(k+D)={a}(k)$ for all $k\geq 0$. $\omega_{N}=e^{\frac{2\pi\sqrt{-1}}{N}}=e^{\frac{2\pi i}{N}}$ is a $N$-th primitive complex root of unity.
Let $\mathbf{a}$ and $\mathbf{b}$ be unimodular sequences with length $D$. Then the aperiodic cross ambiguity function (cross-AF) of $\mathbf{a}$ and $\mathbf{b}$ at time shift $\tau$ and Doppler shift $\nu$ is defined by:
$\widehat{AF}_{\mathbf{a},\mathbf{b}}(\tau,\nu)=\left\{\begin{array}[]{cl}\sum\limits_{t=0}^{D-1-\tau}a(t)\overline{b}(t+\tau)\omega_{D}^{\nu t},&0\leq\tau\leq D-1,\\ \sum\limits_{t=-\tau}^{D-1}a(t)\overline{b}(t+\tau)\omega_{D}^{\nu t},&-(D-1)\leq\tau<0,\\ 0,&|\tau|\geq D.\end{array}\right.$
where for $c\in\mathbb{C}$, $\overline{c}$ represents complex conjugation of $c$.
If both of $\mathbf{a}$ and $\mathbf{b}$ are periodic with period $D$, the periodic cross-AF of $\mathbf{a}$ and $\mathbf{b}$ is defined by:
${AF}_{\mathbf{a},\mathbf{b}}(\tau,\nu)=\sum\limits_{t=0}^{D-1}a(t)\overline{b}(t+\tau)\omega_{D}^{\nu t},$
where $t+\tau$ to be taken modulo the period $D$. We denote $\widehat{AF}_{\mathbf{a}}(\tau,\nu)=\widehat{AF}_{\mathbf{a},\mathbf{a}}(\tau,\nu)$ and ${AF}_{\mathbf{a}}(\tau,\nu)={AF}_{\mathbf{a},\mathbf{a}}(\tau,\nu)$, called the auto-AF of $\mathbf{a}$, aperiodic and periodic, respectively.
For a sequence set $\mathcal{S}$ that consists of $M$ sequences of length $D$, the maximum periodic AF magnitude of $\mathcal{S}$ over a region $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})\subseteq(-D,D)\times(-D,D)$ is defined as
$\theta_{\text{max}}(\mathcal{S})=\{\theta_{\text{A}}(\mathcal{S}),\theta_{\text{C}}(\mathcal{S})\},$
where $\theta_{\text{A}}(\mathcal{S})$ denotes the maximum periodic auto-AF magnitude by
$\theta_{\text{A}}(\mathcal{S})={\text{max}}\{|{AF}_{\mathbf{a}}(\tau,\nu)|:\mathbf{a}\in\mathcal{S},(0,0)\neq(\tau,\nu)\in\Pi\}$
and $\theta_{\text{C}}(\mathcal{S})$ denotes the maximum periodic cross-AF magnitude by
$\theta_{\text{C}}(\mathcal{S})={\text{max}}\{|{AF}_{\mathbf{a},\mathbf{b}}(\tau,\nu)|:\mathbf{a}\neq\mathbf{b}\in\mathcal{S},(\tau,\nu)\in\Pi\}$
Such a sequence set $\mathcal{S}$ is denoted by $(M,D,\Pi,\theta_{\text{max}})$-LAZ periodic sequence set, where $M$ is the set size, $D$ is the length of sequence, $\Pi$ is the low ambiguity zone, as well as $\theta_{\text{max}}$ denotes the maximum periodic AF magnitude in region $\Pi$.
For aperiodic case, we define similarly $\hat{\theta}_{\text{max}}(\mathcal{S})=\{\hat{\theta}_{\text{A}}(\mathcal{S}),\hat{\theta}_{\text{C}}(\mathcal{S})\},$ to be the maximum aperiodic AF magnitude over a region $\Pi$, where
$\hat{\theta}_{\text{A}}(\mathcal{S})={\text{max}}\{|\widehat{AF}_{\mathbf{a}}(\tau,\nu)|:\mathbf{a}\in\mathcal{S},(0,0)\neq(\tau,\nu)\in\Pi\}$
$\hat{\theta}_{\text{C}}(\mathcal{S})={\text{max}}\{|\widehat{AF}_{\mathbf{a},\mathbf{b}}(\tau,\nu)|:\mathbf{a}\neq\mathbf{b}\in\mathcal{S},(\tau,\nu)\in\Pi\}\,\,\,\,\,$
Such a sequence set $\mathcal{S}$ is denoted by $(M,D,\Pi,\hat{\theta}_{\text{max}})$-LAZ aperiodic sequence set.
As mentioned, in [12], authors established lower bounds $\Delta$ and $\hat{\Delta}$ of $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$, respectively. The following theorem is a consequence of [12].
Theorem 1.
([12], or [1] Lemma 1 and 2) For any unimodular sequence set $\mathcal{S}=(M,D,\Pi,\theta_{\text{max}})$ or $\mathcal{S}=(M,D,\Pi,\hat{\theta}_{\text{max}})$, where $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$. We have
(i)  $\theta_{\text{max}}\geq\Delta$, where $\Delta=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}$;
(ii)  $\hat{\theta}_{\text{max}}\geq\hat{\Delta}$, where $\hat{\Delta}=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}$.
Further assume that $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}(\geq 1)$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}(\geq 1)$, $\Delta$ and $\hat{\Delta}$ are YZLFLT bounds, $\rho_{\text{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ are conventionally referred to as optimality factors. The optimality factors $\rho_{\text{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ serve as metrics to evaluate how closely the constructed sequences approach the theoretical lower bounds in both periodic and aperiodic cases, respectively. i.e.
$\bullet$  If ${\rho_{\text{LAZ}}}=1$ or $\hat{\rho}_{\text{LAZ}}$=1, the sequence set $\mathcal{S}$ is called optimal in periodic and aperiodic case, respectively;
$\bullet$  If $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=1$ or $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=1$, the sequence set $\mathcal{S}$ is called asymptotically optimal in periodic and aperiodic case, respectively.
II-B Locally Perfect Nonlinear Functions
Let $K\geq N$ be two positive integers, and $f:$ $\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K}$ be a function, where $\mathbb{Z}_{N}=\mathbb{Z}/N\mathbb{Z}$. For $0<Z_{x}\leq N$ and $0<Z_{y}\leq K$, The function $f$ is called a locally perfect nonlinear functions and denoted by $(N,K,Z_{x},Z_{y})$-LPNF if for any $-Z_{y}<b<Z_{y}$ and $-Z_{x}<a<Z_{x}$, $a\neq 0$, the equation $f(x+a)-f(x)=b$ has at most one solution $x\in\mathbb{Z}_{N}$. In what follows, to avoid confusion, we denote $(N,K,Z_{x},Z_{y})$-LPNF, as $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF.
Remark 1.
If $Z_{y}\geq\lceil\frac{K+1}{2}\rceil$, then $-Z_{y}<b<Z_{y}$ covers any element $b$ in $\mathbb{Z}_{K}$.
The notion of LPNF was introduced in [1] and successfully used to construct LAZ sequence sets by interleaving technique. The result is as follows.
Theorem 2.
([1] Theorem 1) From a $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF we can construct a LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$, $D=NK$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $\theta_{\text{max}}=K$, $\hat{\theta}_{\text{max}}=K+Z_{x}-1$.
Wang and Zhou et al.[1] gave a series of LPNFs and constructed a series of LAZ sequence sets with flexible parameters, which are presented in the above theorem. Some of them are asymptotically optimal in both periodic and aperiodic cases. In next section, we will reveal some new constructions of LPNFs, which provides significant improvements over existing frameworks [1]. And then produce a series of LAZ sequence sets with new parameters that are not covered in Theorem 2. Furthermore, a class of the proposed sequence sets demonstrates asymptotically optimal property in both periodic and aperiodic cases.
III New Construction of LPNFs
For $K\geq 3$, let $\mathbb{Z}_{K}^{\ast}$ denote the multiplication group of all invertible elements of the ring $\mathbb{Z}_{K}$. Then $\mathbb{Z}_{K}^{\ast}$ contains $\varphi(K)$ elements, where $\varphi(\cdot)$ is the Euler’s total function. To establish the desired constructions, we need the following lemma.
Lemma 1.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. There is an element $\alpha$ of order $N$ in $\mathbb{Z}_{K}^{\ast}$.
Proof.
By the Chinese remainder theorem, we have an isomorphism of rings:
$\mathbb{Z}_{K}\cong\mathbb{Z}_{p_{1}^{e_{1}}}\times\mathbb{Z}_{p_{2}^{e_{2}}}\times\cdots\times\mathbb{Z}_{p_{s}^{e_{s}}},\,\,\,\,\,\,x\mapsto(x_{1},\cdots x_{s}),$
where $x_{i}\equiv x\,(\bmod\,p_{i}^{e_{i}})$, $1\leq i\leq s$, which also keeps the isomorphism of groups
$\mathbb{Z}_{K}^{\ast}\cong\mathbb{Z}_{p_{1}^{e_{1}}}^{\ast}\times\mathbb{Z}_{p_{2}^{e_{2}}}^{\ast}\times\cdots\times\mathbb{Z}_{p_{s}^{e_{s}}}^{\ast},$
we identify $x$ with the image $(x\,(\bmod\,p_{1}^{e_{1}}),x\,(\bmod\,p_{2}^{e_{2}})\cdots x\,(\bmod\,p_{i}^{e_{i}}))$.
For every $i$, $1\leq i\leq s$, The group $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}$ is cyclic of order $\varphi(p_{i}^{e_{i}})$, so that $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\langle\pi_{i}\rangle$ with $\pi_{i}$ as a generator. On the other hand, $\pi=(\pi_{1},\pi_{2},\cdots,\pi_{s})$ is an element of $\mathbb{Z}_{K}^{\ast}$ with order $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}$, that is, there exist a cyclic subgroup $\langle\pi\rangle=\{1,\pi,\cdots,\pi^{E-1}\}$ of $\mathbb{Z}_{K}^{\ast}$ with order $E$. Next we write $E=NR$ and put $\alpha=\pi^{R}$, which is an element of $\mathbb{Z}_{K}^{\ast}$ of order $N$. ∎
Construction 1.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. We will define the function $f:\,\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K}$ as $f(x)=\alpha^{x}$, where $\alpha$ is an element of order $N$.
Theorem 3.
Let $f$ be a function defined in Construction 1. Then we have
(i)  the function $f(x)=\alpha^{x}$ is $\langle N,K,Z_{x},Z_{y}\rangle$LPNF, where $Z_{y}=K$ and $Z_{x}=min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$.
(ii)  the function $f(x)=\alpha^{x}$ is $\langle N,K,Z_{x},Z_{y}\rangle$LPNF, where $Z_{x}=N$ and $Z_{y}=p$, $p$ is the smallest prime factor of $K$.
Proof.
Recall the function
$f:\,\,\,\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K},\,\,\,\,\,f(x)=\alpha^{x},$
where $\alpha$ is an element of order $N$.
(i)  For $a\in\mathbb{Z}_{N}\setminus\{0\}$ and $b\in\mathbb{Z}_{K}$, consider the solution $x$ in $\mathbb{Z}_{N}$ of equation
$f(x+a)-f(x)=b,$
which is
$\alpha^{x+a}-\alpha^{x}=\alpha^{x}(\alpha^{a}-1)=b,$ (1)
Note that $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$ means $\alpha^{a}-1$ is invertible in $\mathbb{Z}_{K}$. Furthermore, we can obtain the following equivalence relation :
[…]
$\Longleftrightarrow$ $\alpha^{a}-1\not\equiv 0\,(\bmod\,p_{i})\,\,\,\,\,\,(\mathrm{since}\,x\in\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}\Longleftrightarrow p_{i}\nmid x)\,\,\,\,\,\,(1\leq i\leq s)$
$\Longleftrightarrow$ $\pi^{aR}\not\equiv 1\,(\bmod\,p_{i})\,\,\,\,\,\,(1\leq i\leq s)$
$\Longleftrightarrow$ $p_{i}-1\nmid aR\,\,\,\,\,\,(\mathrm{since\,the\,order\,of}\,\pi\,(mod\,p_{i})\,\mathrm{is}\,p_{i}-1)\,\,\,\,\,\,(1\leq i\leq s)$
$\Longleftrightarrow$ $\frac{p_{i}-1}{\gcd(p_{i}-1,R)}\nmid a\,\,\,\,\,\,(1\leq i\leq s).$
Therefore, from the equivalence relation we conclude that if $-Z_{x}<a<Z_{x}$ and $a\neq 0$, where $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, then $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$ and reformulate the equation (3) as $\alpha^{x}=b(\alpha^{a}-1)$. Hence, if $b(\alpha^{a}-1)\in\langle\alpha\rangle$, the equation (1) must have a unique solution $x\in\mathbb{Z}_{N}$. Conversely, the equation (1) has no solution in $\mathbb{Z}_{N}$.
(ii)  For any integers $-N<a<N$ with $a\neq 0$ and $-p<b<p$, consider the number of solutions to
$f(x+a)-f(x)\equiv b\,(\rm mod\,K)$ (3)
over $x\in\mathbb{Z}_{N}$, where $p$ denotes the smallest prime factor of $K$. Next we distinguish the two cases $b=0$ and $b\neq 0$.
if $b=0$, then we get $\alpha^{x}(\alpha^{a}-1)\equiv 0\,(\rm mod\,K)$. Since $\alpha\in\mathbb{Z}_{K}^{\ast}$, we have $\alpha^{x}\in\mathbb{Z}_{K}^{\ast}$, and thus $\alpha^{a}-1\equiv 0\,(\rm mod\,K)$, i.e., $\alpha^{a}\equiv 1\,(\rm mod\,K)$. Recall that the order of $\alpha$ mod $K$ is $N$, so $N\,|\,a$, which contradicts to $-N<a<N$ with $a\neq 0$. Therefore, the equation (3) has no solution in $\mathbb{Z}_{N}$.
if $b\neq 0$, then we have $\alpha^{x}(\alpha^{a}-1)\equiv b\,(\rm mod\,K)$. Because $p$ is the smallest prime factor of $K$, for $1\leq i\leq s$, all $p_{i}$ hold $p_{i}\geq p$ and $-p_{i}<b<p_{i}$. By combining $b\neq 0$, we get $p_{i}\nmid b$, that is, $b\not\equiv 0\,(\rm mod\,p_{i})$, which yields $b\in\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}$ from $\mathbb{Z}_{p_{i}^{e_{i}}}^{\ast}=\mathbb{Z}_{p_{i}^{e_{i}}}\setminus p_{i}\mathbb{Z}_{p_{i}^{e_{i}}}$.Applying the Chinese remainder theorem, we obtain that $b$ is invertible in $\mathbb{Z}_{K}$, i.e., $b\in\mathbb{Z}_{K}^{\ast}$. Note that $\alpha^{x}\in\mathbb{Z}_{K}^{\ast}$, which leads to $\alpha^{a}-1\equiv\alpha^{-x}b\,(\rm mod\,K)$, and hence, it also implies $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$. Together with $\alpha^{x}=(\alpha^{a}-1)^{-1}b\in\mathbb{Z}_{K}^{\ast}$ and the order of $\alpha$ mod $K$ is $N$, we conclude that equation (3) admits a unique solution $x\in\mathbb{Z}_{N}$.
Consequently, the equation (3) has at most a solution. This completes the proof of Theorem 3. ∎
Example 1.
Let $K=5^{2}=25$, $N=20$, and we choose an element $\alpha=2$ of order $20$ to define the function
$f:\,\,\,\mathbb{Z}_{20}\rightarrow\mathbb{Z}_{25},\,\,\,\,\,f(x)=2^{x}.$
  x 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19
  $f(x)$ 1 2 4 8 16 7 14 3 6 12 24 23 21 17 9 18 11 22 19 13
For $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}=4$, $Z_{y}=K=20$ and take $a\in(-\mathbb{Z}_{4},\mathbb{Z}_{4})\setminus\{0\}$, we have
$\left\{\begin{array}[]{ll}\{f(x-3)-f(x):0\leq x\leq 19\}=\{21,17,9,18,11,22,19,13,1,2,4,8,16,7,14,3,6,12,24,23\}\\ \{f(x-2)-f(x):0\leq x\leq 19\}=\{18,11,22,19,13,1,2,4,8,16,7,14,3,6,12,24,23,21,17,9\}\\ \{f(x-1)-f(x):0\leq x\leq 19\}=\{12,24,23,21,17,9,18,11,22,19,13,1,2,4,8,16,7,14,3,6\}\\ \{f(x+1)-f(x):0\leq x\leq 19\}=\{1,2,4,8,16,7,14,3,6,12,24,23,21,17,9,18,11,22,19,13\}\\ \{f(x+2)-f(x):0\leq x\leq 19\}=\{3,6,12,24,23,21,17,9,18,11,22,19,13,1,2,4,8,16,7,14\}\\ \{f(x+3)-f(x):0\leq x\leq 19\}=\{7,14,3,6,12,24,23,21,17,9,18,11,22,19,13,1,2,4,8,16\}\end{array}\right.$
These indicate that $f(x)=2^{x}$ is a locally perfect nonlinear function, which aligns with our Theorem 3 (i).
Example 2.
Let $K=3^{2}=9$, $N=6$, and we choose an element $\alpha=2$ of order $6$ to define the function
$f:\,\,\,\mathbb{Z}_{6}\rightarrow\mathbb{Z}_{9},\,\,\,\,\,f(x)=2^{x}.$
  x 0 1 2 3 4 5
  $f(x)$ 1 2 4 8 7 5
For $Z_{x}=N=6$, $Z_{y}=p=3$ and take $a\in(-\mathbb{Z}_{6},\mathbb{Z}_{6})\setminus\{0\}$, we have
$\left\{\begin{array}[]{ll}\{f(x-5)-f(x):0\leq x\leq 5\}=\{1,2,4,8,7,5\}\\ \{f(x-4)-f(x):0\leq x\leq 5\}=\{3,6,3,6,3,6\}\\ \{f(x-3)-f(x):0\leq x\leq 5\}=\{7,5,1,2,4,8\}\\ \{f(x-2)-f(x):0\leq x\leq 5\}=\{6,3,6,3,6,3\}\\ \{f(x-1)-f(x):0\leq x\leq 5\}=\{4,8,7,5,1,2\}\\ \{f(x+1)-f(x):0\leq x\leq 5\}=\{1,2,4,8,7,5\}\\ \{f(x+2)-f(x):0\leq x\leq 5\}=\{3,6,3,6,3,6\}\\ \{f(x+3)-f(x):0\leq x\leq 5\}=\{7,5,1,2,4,8\}\\ \{f(x+4)-f(x):0\leq x\leq 5\}=\{6,3,6,3,6,3\}\\ \{f(x+5)-f(x):0\leq x\leq 5\}=\{4,8,7,5,1,2\}\end{array}\right.$
We find that the solutions of $f(x+4)-f(x)=3$ are $1,3,5$. Moreover, $f(x-4)-f(x)=3$, $f(x-2)-f(x)=3$, and $f(x+2)-f(x)=3$ have more than one solution. These imply that $f(x)=2^{x}$ is a locally perfect nonlinear function if $-6<a<6$, $a\neq 0$ and $-3<b<3$, which aligns with our Theorem 3 (ii).
The next characterizes a construction of LPNFs with new parameters, which are employed to construct LAZ sequence sets.
Construction 2.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, the $p_{i}$ are distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$. Denote $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}})=p_{i}^{e_{i}-1}(p_{i}-1):1\leq i\leq s\}=NR$ with $N\geq 2$, $N,R\in\mathbb{Z}$. Put $K_{1}\geq K$, we will define the function $f:\,\mathbb{Z}_{N}\rightarrow\mathbb{Z}_{K_{1}}$ as $f(x)=\langle\alpha^{x}\rangle_{K}$, where $\alpha$ is an element of order $N$, $\langle\alpha\rangle_{K}$ is the least positive integer of $\alpha$ modulo $K$.
Theorem 4.
Let $f$ be a function defined in Construction 2. Then we have $f$ is $\langle N,K,Z_{x},Z_{y}\rangle$-LPNF, where $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $Z_{y}=K_{1}-K+2$.
Proof.
By the definition of LPNFs, it suffices to consider the equation
$f(x+a)-f(x)=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}\equiv b\,(\rm mod\,K_{1}),$ (4)
which has at most one solution $x\in\mathbb{Z}_{N}$ when $-Z_{x}<a<Z_{x}$ with $a\neq 0$ and $-Z_{y}<b<Z_{y}$.
Since $11\leq f(x),f(x+a)\leq K-$ 1, we must have
$-(K-2)\leq f(x+a)-f(x)\leq K-2.$
Thus, from $-(K_{1}-K+2)<b<K_{1}-K+2$ yields
$-(K_{1}-1)\leq f(x+a)-f(x)-b\leq K_{1}-1.$
Furthermore, the equation (4) implies $K_{1}\,|\,f(x+a)-f(x)-b$, and so, it denotes the conventional arithmetic operations applied to integers $\mathbb{Z}$, i.e.,
$f(x+a)-f(x)-b=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}-b=0.$ (5)
Consequently, the equation (5) is valid in $\mathbb{Z}_{K}$, that is,
$f(x+a)-f(x)=\langle\alpha^{x+a}\rangle_{K}-\langle\alpha^{x}\rangle_{K}\equiv b\,(\rm mod\,K).$
But $f(x)=\langle\alpha^{x}\rangle_{K}\equiv\alpha^{x}\,(\rm mod\,K)$, so that we can get
$\alpha^{x+a}-\alpha^{x}\equiv b\,(\rm mod\,K).$ (6)
For $-{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}<a<{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ with $a\neq 0$, applying Theorem 3 we can immediately conclude that $\alpha^{a}-1\in\mathbb{Z}_{K}^{\ast}$. Hence, equation (6) has at most a solution in $\mathbb{Z}_{N}$ if $-{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}<a<{\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$ and $-(K_{1}-K+1)<b<K_{1}-K+1$.
Consequently, The equation (4) admits at most one solution. ∎
Example 3.
Let $K=7$, $K_{1}=10$, $N=6$, and we choose an element $\alpha=5$ of order $6$ to define the function
$f:\,\,\,\mathbb{Z}_{6}\rightarrow\mathbb{Z}_{10},\,\,\,\,\,f(x)=\langle 5^{x}\rangle_{7}.$
  x 0 1 2 3 4 5
  $f(x)$ 1 5 4 6 2 3
For $Z_{x}={\text{min}}\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}=6$, $Z_{y}=K_{1}-K+2=5$ and take $a\in(-\mathbb{Z}_{6},\mathbb{Z}_{6})\setminus\{0\}$, we have
$\left\{\begin{array}[]{ll}\{f(x-5)-f(x):0\leq x\leq 5\}=\{4,9,2,6,1,8\}\\ \{f(x-4)-f(x):0\leq x\leq 5\}=\{3,1,8,7,9,2\}\\ \{f(x-3)-f(x):0\leq x\leq 5\}=\{5,7,9,5,3,1\}\\ \{f(x-2)-f(x):0\leq x\leq 5\}=\{1,8,7,9,2,3\}\\ \{f(x-1)-f(x):0\leq x\leq 5\}=\{2,6,1,8,4,9\}\\ \{f(x+1)-f(x):0\leq x\leq 5\}=\{4,9,2,6,1,8\}\\ \{f(x+2)-f(x):0\leq x\leq 5\}=\{3,1,8,7,9,2\}\\ \{f(x+3)-f(x):0\leq x\leq 5\}=\{5,7,9,5,3,1\}\\ \{f(x+4)-f(x):0\leq x\leq 5\}=\{1,8,7,9,2,3\}\\ \{f(x+5)-f(x):0\leq x\leq 5\}=\{2,6,1,8,4,9\}\end{array}\right.$
These show that the solutions of $f(x+3)-f(x)=5$ and $f(x-3)-f(x)=5$ are $0,3$. we conclude that $f(x)=2^{x}$ is a locally perfect nonlinear function if $-6<a<6$, $a\neq 0$ and $-5<b<5$, which aligns with our Theorem 4.
IV New LAZ sequence sets
By applying the locally perfect nonlinear functions (LPNFs) in Construction 1 and 2 to the above Theorem 2, we directly get the following LAZ sequence sets.
Theorem 5.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. Then
(i)  From Construction 1 (i), we have a LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$ is the number of sequences, $D=NK$ is the length of sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K$, $Z_{x}=min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, $\theta_{\text{max}}=K$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1$.
(ii)  From Construction 1 (ii), we have a LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$ is the number of sequences, $D=NK$ is the length of sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{x}=N$, $Z_{y}=p$, $p$ is the smallest prime factor of $K$, $\theta_{\text{max}}=K$, and $\hat{\theta}_{\text{max}}=K+N-1$.
Theorem 6.
Let $K=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{s}^{e_{s}}$, where $s\geq 1$, $\{p_{1},p_{2},\cdots,p_{s}\}$ be distinct odd primes and $e_{i}\geq 1$ with $1\leq i\leq s$, $E=\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}=NR$ with $N\geq 2$. Then, from Construction 2, we have a LAZ sequence set $\mathcal{S}$ with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N$ is the number of sequences, $D=NK_{1}$ is the length of sequences, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K_{1}-K+1$, $K\leq K_{1}$, $Z_{x}=min\{\frac{p_{i}-1}{\gcd(p_{i}-1,R)}:1\leq i\leq s\}$, $\theta_{\text{max}}=K_{1}$, and $\hat{\theta}_{\text{max}}=K_{1}+Z_{x}-1$.
For the $(N,NK,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$-LAZ sequence set constructed in [1], the value of the parameter $N$ is only odd. But for the construction in Theorem 3, the value of parameter $N$, as a factor of $\textrm{lcm}\{\varphi(p_{i}^{e_{i}}):1\leq i\leq s\}$, can be chosen to be even. Thus the LAZ sequence set in Theorem 5 has more general and flexible parameters, and then one series of them is asymptotically optimal as shown in following corollary.
Corollary 1.
Let $K=p^{e}$, where $p$ is an odd prime and $e\geq 1$, and let $N=\varphi(p^{e})=p^{e-1}(p-1)$. There exist LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N=\varphi(p^{e})$, $D=NK=p^{e}\varphi(p^{e})$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$, $\theta_{\text{max}}=K=p^{e}$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p^{e}+p-2$. Moreover, the series of LAZ sequence set is asymptotically optimal in periodic case when $p\rightarrow\infty$, and asymptotically optimal in aperiodic case when $e\geq 2$ and $p\rightarrow\infty$.
Proof.
The parameters $M$, $D$, $Z_{x}$, $Z_{y}$, $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$ can be seen directly from Theorem 3 and 5 by taking $s=R=1$. By Theorem 1, the YZLFLT bounds $\Delta$ and $\hat{\Delta}$ can be derived as follows:
$\left\{\begin{array}[]{ll}\Delta=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}=N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}},\\ \hat{\Delta}=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}=N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}},\end{array}\right.$
thus, we have
$\frac{\Delta}{\theta_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}}}{K}$
$=$ $\frac{N\sqrt{\frac{p-2}{N(p-1)-1}}}{\sqrt{K}}.$
Observe that if $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\Delta}{\theta_{\text{max}}}=\lim\limits_{p\to\infty}\sqrt{\frac{N}{K}}=1$ since $\frac{K}{N}=\frac{p^{e}}{p^{e-1}(p-1)}=\frac{p}{p-1}\rightarrow 1$. Furthermore, we have
$\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}}}{K+p-2}$
$=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-2)-(p-2)}{NKN(p-1)+N(p-2)(p-1)-(p+NK-2)}}}{K+p-2}.$
Observe that if $e\geq 2$ and $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}=\lim\limits_{p\to\infty}\frac{N\sqrt{K}}{K}\sqrt{\frac{NK(p-2)}{NKN(p-1)}}=\lim\limits_{p\to\infty}\sqrt{\frac{N}{K}}=1$ since $\hat{\theta}_{\text{max}}=p^{e}+p-2\rightarrow K$ and $\frac{K}{N}\rightarrow 1$. The proof of Corollary 1 is completed. ∎
Now we derive the case of twin-prime.
Corollary 2.
Let $K=p(p+2)$, where $p$ and $p+2$ are odd prime, and let $N=\textrm{lcm}(p-1,p+1)=\frac{p^{2}-1}{2}$. There exist LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N=\frac{p^{2}-1}{2}$, $D=NK=\frac{p(p+2)(p^{2}-1)}{2}$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p^{e}$, $Z_{x}=p-1$, $\theta_{\text{max}}=K=p(p+2)$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p(p+2)+p-2$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$ and $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
Proof.
The parameters $M$, $D$, $Z_{x}$, $Z_{y}$, $\theta_{\text{max}}$ and $\hat{\theta}_{\text{max}}$ can be seen directly from Theorem 3 and 5 by taking $s=2$ and $R=1$. By Theorem 1, the YZLFLT bounds $\Delta$ and $\hat{\Delta}$ can be derived as follows:
$\left\{\begin{array}[]{ll}\Delta=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D}{D(MZ_{x}-1)}}=N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}},\\ \hat{\Delta}=\frac{D}{\sqrt{Z_{y}}}\sqrt{\frac{MZ_{x}Z_{y}-D-Z_{x}+1}{(MZ_{x}-1)(D+Z_{x}-1)}}=N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}},\end{array}\right.$
thus, we have
$\frac{\Delta}{\theta_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{p-2}{N(p-1)-1}}}{K}$
$=$ $\frac{N\sqrt{\frac{p-2}{N(p-1)-1}}}{\sqrt{K}}.$
Observe that if $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\theta_{\text{max}}}{\Delta}=\lim\limits_{p\to\infty}\sqrt{\frac{K}{N}}=\sqrt{2}$ since $\frac{K}{N}=\frac{p(p+2)}{\frac{p^{2}-1}{2}}=\frac{2p^{2}+4p}{p^{2}-1}\rightarrow 2$. Furthermore, we have
$\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}$ $=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-1)-(NK+p-2)}{(NK+p-2)(N(p-1)-1)}}}{K+p-2}$
$=$ $\frac{N\sqrt{K}\sqrt{\frac{NK(p-2)-(p-2)}{NKN(p-1)+N(p-2)(p-1)-(p+NK-2)}}}{K+p-2}.$
Observe that if $p\rightarrow\infty$, then $\lim\limits_{p\to\infty}\frac{\hat{\Delta}}{\hat{\theta}_{\text{max}}}=\lim\limits_{p\to\infty}\frac{N\sqrt{K}}{K}\sqrt{\frac{NK(p-2)}{NKN(p-1)}}=\lim\limits_{p\to\infty}\sqrt{\frac{N}{K}}$ and $\lim\limits_{p\to\infty}\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}=\lim\limits_{p\to\infty}\sqrt{\frac{K}{N}}=\sqrt{2}$ since $\hat{\theta}_{\text{max}}=p(p+2)+p-2\rightarrow K$ and $\frac{K}{N}\rightarrow 2$. The proof of Corollary 2 is completed. ∎
Similarly, we can derive the following corollary, which addresses the case where $K$ is the product of distinct odd primes.
Corollary 3.
Let $K=p_{1}p_{2}\cdots p_{s}$, where $p_{1}<p_{2}<\cdots<p_{s}$ are odd primes, and when $N=E=\textrm{lcm}\{\varphi(p_{i})=p_{i}-1:1\leq i\leq s\}=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)}{2}$. There exist LAZ sequence set with parameters $(M,D,\Pi,\theta_{\text{max}}\,\,or\,\,\hat{\theta}_{\text{max}})$, where $M=N=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)}{2}$, $D=NK=\frac{(p_{1}-1)(p_{2}-1)\cdots(p_{s}-1)p_{1}p_{2}\cdots p_{s}}{2}$, $\Pi=(-Z_{x},Z_{x})\times(-Z_{y},Z_{y})$, $Z_{y}=K=p_{1}p_{2}\cdots p_{s}$, $Z_{x}=p_{1}-1$, $\theta_{\text{max}}=K=p_{1}p_{2}\cdots p_{s}$, and $\hat{\theta}_{\text{max}}=K+Z_{x}-1=p_{1}p_{2}\cdots p_{s}+p_{1}-2$. Moreover, $\lim\limits_{D\to\infty}\rho_{\text{LAZ}}=\sqrt{2}$ and $\lim\limits_{D\to\infty}\hat{\rho}_{\text{LAZ}}=\sqrt{2}$, where $\rho_{\text{LAZ}}=\frac{\theta_{\text{max}}}{\Delta}$ and $\hat{\rho}_{\text{LAZ}}=\frac{\hat{\theta}_{\text{max}}}{\hat{\Delta}}$.
Proof.
The method is the same as in Corollary 1 and 2, and thus it is omitted. ∎
Here, we provide an experimental example demonstrating that values of $N$ can be identified for which the sequence set $\mathcal{S}$ is a LAZ sequence set, achieves asymptotic optimality. Thus, with careful selection of parameters, our sequences are strong candidates for ISAC system.
IV-A Example of Periodic and Aperiodic LAZ sequence set
In this subsection, we present two examples to illustrate our proposed LAZ sequence sets in Theorems 5 and 6.
Example 4.
Let $K=25$, $N=20$, $\alpha=2$ be the element in $\mathbb{Z}_{K}$ of order $20$. we define the function
$f:\,\,\,\mathbb{Z}_{20}\rightarrow\mathbb{Z}_{25},\,\,\,\,\,f(x)=2^{x}.$
Then
  x 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19
  $f(x)$ 1 2 4 8 16 7 14 3 6 12 24 23 21 17 9 18 11 22 19 13
The sequence set $\mathcal{S}$ is created using Theorem 5. It can be demonstrated that $\mathcal{S}$ forms a set of periodic sequences $(20,500,\Pi_{1},25)$-LAZ and $(20,500,\Pi_{2},25)$-LAZ, where $\Pi_{1}=(-4,4)\times(-25,25)$ and $\Pi_{2}=(-20,20)\times(-5,5)$. A glimpse of the periodic auto-AF and cross-AF of the sequences in $\mathcal{S}$ can be seen in Fig. 1.
Figure 1: A glimpse of the periodic auto-AF and cross-AF of the sequence set $\mathcal{S}$ in Example 4.
Furthermore, it can be shown that $\mathcal{S}$ also constitutes an aperiodic set of $(20,500,\Pi_{1},30)$-LAZ and $(20,500,\Pi_{2},30)$-LAZ sequences, where $\Pi_{1}=(-4,4)\times(-25,25)$ and $\Pi_{2}=(-20,20)\times(-5,5)$. A glimpse of the aperiodic auto-AF and cross-AF of the sequences in $\mathcal{S}$ can be seen in Fig. 2.
Figure 2: A glimpse of the aperiodic auto-AF and cross-AF of the sequence set $\mathcal{S}$ in Example 4.
Example 5.
Let $K=7$, $K_{1}=10$, $N=6$, $\alpha=5$ be the element in $\mathbb{Z}_{K}$ of order $6$. we define the function
$f:\,\,\,\mathbb{Z}_{6}\rightarrow\mathbb{Z}_{10},\,\,\,\,\,f(x)=\langle 5^{x}\rangle_{K}.$
Regard $f(x)$ as a function in $\mathbb{Z}_{K_{1}}$. Then
  x 0 1 2 3 4 5
  $f(x)$ 1 5 4 6 2 3
The sequence set $\mathcal{S}$ is created using Theorem 6. It can be demonstrated that $\mathcal{S}$ forms a periodic $(6,60,\Pi,10)$-LAZ sequence set, where $\Pi=(-6,6)\times(-5,5)$. A glimpse of the periodic auto-AF and cross-AF of the sequences in $\mathcal{S}$ can be seen in Fig. 3.
Figure 3: A glimpse of the periodic auto-AF and cross-AF of the sequence set $\mathcal{S}$ in Example 5.
Additionally, it can be shown that $\mathcal{S}$ also constitutes an aperiodic $(6,60,\Pi,15)$-LAZ sequence set, where $\Pi=(-6,6)\times(-5,5)$. A glimpse of the aperiodic auto-AF and cross-AF of the sequences in $\mathcal{S}$ can be seen in Fig. 4.
Figure 4: A glimpse of the aperiodic auto-AF and cross-AF of the sequence set $\mathcal{S}$ in Example 5.
IV-B Parameters of the Derived Periodic and Aperiodic LAZ Sequence Sets
In Tables II, III, and IV, we present several parameters of the proposed periodic LAZ sequence sets in Theorems 5 and 6, along with their corresponding optimality factor, $\rho_{\text{LAZ}}.$ Furthermore, we provide parameters of the proposed aperiodic LAZ sequence sets in Theorems 5 and 6, along with their corresponding optimality factor, $\hat{\rho}_{\text{LAZ}}$, in Tables V and VI, respectively. It is obvious from these tables that the optimality factors ${\rho}_{\mathrm{LAZ}}$ and $\hat{\rho}_{\text{LAZ}}$ of the constructed LAZ sequence sets are asymptotically approaching $1$. Thus, with careful selection of parameters, our sequences are strong candidates for ISAC system.
TABLE II: Parameters of the proposed periodic LAZ sequence set in Theorem 5(i)
Set size Length Low ambiguity zone Maximum periodic AF Optimality factor
$N$ $NK$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
6 54 $(-3,3)\times(-9,9)$ 9 1.4577
20 500 $(-5,5)\times(-25,25)$ 25 1.2437
42 2058 $(-7,7)\times(-49,49)$ 49 1.1647
110 13310 $(-11,11)\times(-121,121)$ 121 1.0995
156 26364 $(-13,13)\times(-169,169)$ 169 1.0831
272 78608 $(-17,17)\times(-289,289)$ 289 1.0624
506 267674 $(-23,23)\times(-529,529)$ 529 1.0454
812 682892 $(-29,29)\times(-841,841)$ 841 1.0357
930 893730 $(-31,31)\times(-961,961)$ 961 1.0333
1332 1823508 $(-37,37)\times(-1369,1369)$ 1369 1.0278
TABLE III: Parameters of the proposed periodic LAZ sequence set in Theorem 5(ii)
Set size Length Low ambiguity zone Maximum periodic AF Optimality factor
$N$ $NK$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
6 54 $(-6,6)\times(-3,3)$ 9 1.7078
20 500 $(-20,20)\times(-5,5)$ 25 1.2894
42 2058 $(-42,42)\times(-7,7)$ 49 1.1829
110 13310 $(-110,110)\times(-11,11)$ 121 1.1055
156 26364 $(-156,156)\times(-13,13)$ 169 1.0871
272 78608 $(-272,272)\times(-17,17)$ 289 1.0646
506 267674 $(-506,506)\times(-23,23)$ 529 1.0465
812 682892 $(-812,812)\times(-29,29)$ 841 1.0364
930 893730 $(-930,930)\times(-31,31)$ 961 1.0339
1332 1823508 $(-1332,1332)\times(-37,37)$ 1369 1.0282
TABLE IV: Parameters of the proposed periodic LAZ sequence set in Theorem 6
Set size Length Low ambiguity zone Maximum periodic AF Optimality factor
$N$ $NK_{1}$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
6 54 $(-6,6)\times(-4,4)$ 9 1.5275
12 204 $(-12,12)\times(-6,6)$ 17 1.3571
22 638 $(-22,22)\times(-8,8)$ 29 1.2550
36 1584 $(-36,36)\times(-9,9)$ 44 1.1888
52 3224 $(-52,52)\times(-11,11)$ 62 1.1562
66 5082 $(-66,66)\times(-12,12)$ 77 1.1367
78 7020 $(-78,78)\times(-13,13)$ 90 1.1252
96 10464 $(-96,96)\times(-14,14)$ 109 1.1115
108 13176 $(-108,108)\times(-15,15)$ 122 1.1052
126 17766 $(-126,126)\times(-16,16)$ 141 1.0969
TABLE V: Parameters of the proposed aperiodic LAZ sequence set Theorem 5(i)
Set size Length Low ambiguity zone Maximum periodic AF Optimality factor
$N$ $NK$ $\Pi$ magnitude $\theta_{\max}$ $\rho_{\mathrm{LAZ}}$
6 54 $(-3,3)\times(-9,9)$ 11 1.8314
20 500 $(-5,5)\times(-25,25)$ 29 1.4499
42 2058 $(-7,7)\times(-49,49)$ 55 1.3095
110 13310 $(-11,11)\times(-121,121)$ 131 1.1909
156 26364 $(-13,13)\times(-169,169)$ 181 1.1603
272 78608 $(-17,17)\times(-289,289)$ 305 1.1213
506 267674 $(-23,23)\times(-529,529)$ 551 1.0889
812 682892 $(-29,29)\times(-841,841)$ 869 1.0702
930 893730 $(-31,31)\times(-961,961)$ 991 1.0656
1332 1823508 $(-37,37)\times(-1369,1369)$ 1405 1.0548
TABLE VI: Parameters of the proposed aperiodic LAZ sequence set Theorem 6
Set size Length Low ambiguity zone Maximum aperiodic AF Optimality factor
$N$ $NK_{1}$ $\Pi$ magnitude $\hat{\theta}_{\max}$ $\hat{\rho}_{\mathrm{LAZ}}$
42 3990 $(-6,6)\times(-48,48)$ 100 1.9319
156 62244 $(-12,12)\times(-232,232)$ 410 1.7752
506 445786 $(-22,22)\times(-354,354)$ 902 1.4345
1332 2521476 $(-36,36)\times(-526,526)$ 1928 1.2798
2756 9725924 $(-52,52)\times(-722,722)$ 3580 1.2060
4422 23790360 $(-66,66)\times(-893,893)$ 5445 1.1711
6162 44853198 $(-78,78)\times(-1040,1040)$ 7356 1.1512
9312 99340416 $(-96,96)\times(-1261,1261)$ 10763 1.1308
11772 156402792 $(-108,108)\times(-1407,1407)$ 13393 1.1210
16002 284275530 $(-126,126)\times(-1638,1638)$ 17890 1.1099
V Concluding Remarks
In this paper, we introduce some constructions of Locally Perfect Nonlinear Functions (LPNFs) with new parameters. By leveraging their structural features, we further propose a series of LAZ sequence sets with more flexible parameters. Notably, one class of these sequence sets achieves both periodic and aperiodic asymptotic optimality.